40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2047     Entry time: Sat Oct 3 14:53:24 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: more progress 

Late last night after the ETMY settled down from the quake I made some more progress in locking, with the handoff to RF CARM succeeding once.  The final reduction of the CARM offset to zero didn't work, however.

ELOG V3.1.3-